OpenStudy (anonymous):

7/9x+7/45=1/5x-1/15. Please show how to solve and explain steps to solve.

OpenStudy (anonymous):

Can anyone help with this problem?

OpenStudy (akashdeepdeb):

\(\frac{7x}{9} + \frac{7}{45} = \frac{1x}{5} - \frac{1}{15}\) Is this the question? :)

OpenStudy (anonymous):

The "x"'s in both of the fractions go to the right. Like this: \[\frac{ 7 }{ 9 }x+\frac{ 7 }{ 45 }=\frac{ 1 }{ 5 }x-\frac{ 1 }{ 15 }\]

OpenStudy (akashdeepdeb):

Hehe...It means the same thing! :D

OpenStudy (anonymous):

Oh ok. Can you show me how to solve it?

OpenStudy (akashdeepdeb):

Sure! :) Take all the terms consisting x onto the other side like this: \[\frac{ 7x }{ 9 } - \frac{ x }{ 5 } = \frac{ -7 }{ 45 } - \frac{ 1 }{ 15 }\] NOW TAKE LCM ON BOTH SIDES [45] \[\frac{ 35x - 9x }{ 45 }= \frac{ -7 - 3 }{ 45 }\] MULTIPLY BY 45 \[35x - 9x = -10\] SOLVE SIMPLY 26x = -10 x = \(\frac{-10}{26}\) SIMPLIFY NOW! :) Understood? :)

OpenStudy (anonymous):

Only thing I don't understand is the part of -7-3. How did you get -7 and -3?

OpenStudy (anonymous):

Nevermind. I figured it out.

OpenStudy (akashdeepdeb):

Did you understand the question then?

OpenStudy (anonymous):

Yes. Thank you.

OpenStudy (akashdeepdeb):

:)